# From Symmetry to Structure: Gauge-Invariant Operators in Multi-Matrix Quantum Mechanics

### v1

@@ 1 Introduction
The number of primary invariants is given by the Krull dimension of the algebra of gauge invariants. For the multi-matrix quantum mechanics of $d$ matrices the Krull dimension is given by $1+(d-1)N^{2}$. In Section 2 we explain how the number of primary invariants can be computed by performing a complete gauge fixing of the theory and counting the number of degrees of freedom that remain. We explain how to reproduce the Krull dimension for multi-matrix models and then go onto to predict the number of primary invariants in a theory of bifundamental fields. In a theory with $f$ bifundamental field we find that the number of primary invariants is given by
@@ 2 Counting Primary Invariants
[…]
@@ 2.1 Multiple Matrices
Notice that all the induced phases depend only on differences $\phi_{a}-\phi_{b}$. Consequently, a common shift $\phi_{a}\to\phi_{a}+\alpha$ leaves all matrices invariant. Hence, of the $N$ residual phases $\phi_{a}$, only $N-1$ are physically meaningful and can be used to further reduce the degrees of freedom.
Since $X^{2}$ is Hermitian, its matrix elements satisfy $m_{ab}=m_{ba}^{*}$. The off-diagonal entries thus contain non-trivial phases. We can use the $N-1$ phase differences $\phi_{a}-\phi_{a+1}$, for $a=1,2,\dots,N-1$, to eliminate the phases of $m_{a\,a+1}$, setting each to be real. These are the additional fields we can fix using the residual symmetry.
We now count the remaining degrees of freedom. There are $(d-1)N^{2}$ matrix elements in the fields $X^{a}$ for $a=2,\dots,d$, and $N$ eigenvalues in the diagonal matrix $X^{1}$. After fixing $N-1$ phases in $X^{2}$ as described above, we find
@@
independent parameters. This correctly counts the number of primary invariants derived from a system of $d>1$ $N\times N$ Hermitian matrices, and agrees with computation of the Krull dimension performed in herstein .
@@
matching the number of eigenvales.
@@ 2.2 Bifundamental fields
[…]
@@
that preserve the diagonal and real form of $D$. These $N$ phases are subject to an overall $U(1)$ redundancy (a common shift), so we can use $N-1$ residual parameters to further simplify the remaining fields.
In particular, we apply these residual transformations to $\Gamma^{2}$, which remains unfixed. Since $\Gamma^{2}$ is a complex matrix, the phases in its off-diagonal entries are complex. As in the case of a single Hermitian matrix, we can use $N-1$ relative phase rotations to eliminate the imaginary parts of $N-1$ off-diagonal entries. For example, we may choose to set the phases of the entries $(\Gamma^{2})_{a,a+1}$ to zero for $a=1,\dots,N-1$.
@@ 3 Counting of Secondary Invariants
$H$ $=$ 12Tr (${1\over 2}\mbox{Tr\,}(\Pi^{1}\Pi^{1})+{1\over 2}\mbox{Tr\,}(X^{1}X^{1})+{1\over 2}\mbox{Tr\,}(\Pi^{2}\Pi^{2})+{1\over 2}\mbox{Tr\,}(X^{2}X^{2})$ (20)
@@
The eigenstates are given by polynomials in the creation operators of a fixed degree. We can refine this further by grading with respect to degree in $A^{1\,\dagger}$ and degree in $A^{2\,\dagger}$ separately. Any multitrace operator, constructed from $n$ $A^{1\,\dagger}$ fields and $m$ $A^{2\,\dagger}$ fields can be written as
@@
where $\delta_{[\sigma]_{r}[\tau]_{r}}$ is 1 if $[\sigma]_{r}$ and $[\tau]_{r}$ belong to the same restricted conjugacy class and it is zero otherwise. If two permutations $\rho_{1}$ and $\rho_{2}$ are related as
@@
We have indicated where our interest is: in the coefficient of the term $x^{\alpha N^{2}}$ which counts the number of operators of degree $\alpha N^{2}$. We have in mind that $\alpha$ is held fixed as we take $N$ large. This formula is useful because the asymptotic behaviour of the Littlewood-Richardson number needed to estimate the behaviour of the coefficient of $x^{q}$ with $q$ of order $N^{2}$ is known Stanley1 ; Stanley2 ; PakPanova . If $R$ has $n$ boxes, $r$ has $k$ boxes and $s$ has $n-k$ boxes, when $n$, $k$ and $n-k$ are all large, the largest Littlewood-Richardson coefficient is given by
@@
The result (33) maybe disturbed in our setting, by taking finite $N$ effects into account. This implies that the Young diagram $R$ must not have more than $N$ rows, It is therefore necessary to inquire about the shape of the Young diagrams $R$, $r$ and $s$ that attain the maximum Littleoowd-Richardson numbers. It turns out, that all three Young diagrams must take the Vershik-Kerov-Logan-Shepp (VKLS) shapeVK1 ; VK2 ; LS . These Young diagrams have a triangular shape and are not cut off by finite $N$ effects for Young diagrams with $\alpha N^{2}$ boxes, as long as $\alpha$ is not too large. Young diagrams with even more boxes take on the truncated VKLS shape LS .
@@
$d_{\alpha N^{2}}<\hat{d}_{\alpha N^{2}}$ $=$ $e^{((\alpha+1)\log(\alpha+1)-\alpha\log(\alpha))N^{2}}\,x^{\alpha N^{2}}$ (48)
@@
$\alpha\log 2<{1\over 15}\left((15+\alpha)\log(1+{\alpha\over 15})-\alpha\log{\alpha\over 15}\right)$ (55)
so that even the leading growth of the coefficient $d_{\alpha N^{2}}$ can not be reproduced from the contribution of $Z_{p}(x)$. It must be that multiplication of $\hat{Z}_{p}(x)$ by $(1+\sum_{i}c^{s}_{i}x^{i})$ enhances the growth, at least to $e^{\alpha\log(2)N^{2}}$, and this can only happen if $(1+\sum_{i}c^{s}_{i}x^{i})$ has $O(e^{cN^{2}})$ terms with $c$ an $O(1)$ number. This proves that the number of secondary invariants must grow as
$1+\sum_{i}c^{s}_{i}x^{i}$ $=$ $e^{cN^{2}}$ (56)
with $c$ a constant, which is the result we wanted to prove.
@@ Acknowledgements
The work of RdMK is supported by a start up research fund of Huzhou University, a Zhejiang Province talent award and by a Changjiang Scholar award. The work of MK is supported by NRF grants 2023R1A2C200536011, 2020R1A6A1A03047877 (CQUeST). HJRvZ would like to acknowledge support from the “Quantum Technologies for Sustainable Development” grant from the National Institute for Theoretical and Computational Sciences of South Africa (NITHECS).
@@ Appendix A Molien-Weyl Functions
×#($\qquad\times\#(\{n_{A_{1}},n_{A_{2}},p_{I_{1}},p_{I_{2}}\})$
where #($\#(\{n_{A_{1}},n_{A_{2}},p_{I_{1}},p_{I_{2}}\})$ is the number of singlets that can be constructed using $n_{A_{1}}$ $\Gamma^{A_{1}}_{12}$ bifundamental fields, $n_{A_{2}}$ $\Gamma^{A_{2}}_{12}$ bifundamental fields, $p_{I_{1}}$ adjoints of the first $U(N)$ factor and $p_{i_{2}}$ adjoints of the second $U(N)$ factor. Using characters we can write the number of singlets as an integral over $U(N)\times U(N)$ to obtain
@@
×$\times\,\sum_{n_{A_{1}}=0}^{\infty}(w^{A_{2}}_{21})^{n_{A_{2}}}\chi_{{\rm sym}^{n_{A_{2}}}(F)}(U_{2})\chi_{{\rm sym}^{n_{A_{2}}}(\bar{F})}(U_{1})\prod_{I_{1}=1}^{b_{1}}\sum_{p_{I_{1}}=0}^{\infty}(z_{1}^{I_{1}})^{p_{I_{1}}}\chi_{{\rm sym}^{p_{I_{1}}}(adj)}(U_{1})$
×$\qquad\times\prod_{I_{2}=1}^{b_{2}}\sum_{p_{I_{2}}=0}^{\infty}(z_{2}^{I_{2}})^{p_{I_{2}}}\chi_{{\rm sym}^{p_{I_{2}}}(adj)}(U_{2})$
$adj$ stands for the adjoint representation. As always, symmetric products of representations appear above because the fields we are using are bosonic. A useful identity is
@@
×$\times{\rm exp}\left(\sum_{A_{1}=1}^{f}\sum_{m=1}^{\infty}{(w_{12}^{A_{1}})^{m}\over m}\chi_{F}(U_{1}^{m})\chi_{\bar{F}}(U_{2}^{m})+\sum_{A_{2}=1}^{f}\sum_{m=1}^{\infty}{(w_{21}^{A_{2}})^{m}\over m}\chi_{F}(U_{2}^{m})\chi_{\bar{F}}(U_{1}^{m})\right)$
@@
×$\times\prod_{I_{1}=1}^{b_{1}}{1\over(1-z_{1}^{I_{1}})^{N}}\prod_{1\leq k<r\leq N}{1\over(1-{\varepsilon_{1r}\over\varepsilon_{1k}}z_{1}^{I_{1}})(1-{\varepsilon_{1k}\over\varepsilon_{1r}}z_{1}^{I_{1}})}$
×$\times\prod_{I_{2}=1}^{b_{2}}{1\over(1-z_{2}^{I_{2}})^{N}}\prod_{1\leq s<t\leq N}{1\over(1-{\varepsilon_{2t}\over\varepsilon_{2s}}z_{2}^{I_{2}})(1-{\varepsilon_{2s}\over\varepsilon_{2t}}z_{2}^{I_{2}})}$
×$\times\prod_{A_{1}=1}^{f}\prod_{1\leq i_{1},j_{1}\leq N}{1\over(1-w_{12}^{A_{1}}{\varepsilon_{1i_{1}}\over\varepsilon_{2j_{1}}})}\prod_{A_{2}=1}^{f}\prod_{1\leq i_{2},j_{2}\leq N}{1\over(1-w_{21}^{A_{2}}{\varepsilon_{2i_{2}}\over\varepsilon_{1j_{2}}})}$
We now simplify the integrand of the above formula. According to (80) and (81) each of the the products $\Delta\bar{\Delta}$ involves a sum over permutations $\sigma$ and $\tau$. For each distinct term in the sum over $\sigma$, consider the distinct change of variables given by a permutation of the eigenvalues $\varepsilon_{1,i}\to\varepsilon_{1,\sigma(i)}$, $\varepsilon_{2,i}\to\varepsilon_{2,\sigma(i)}$. All factors in the integrand are invariant under this change except for the Van der Monde determinants which are both either invariant or both pick up a minus sign. Consequently each product $\Delta\bar{\Delta}$ is invariant. The net result of this change of variables is that the sum over $\sigma$ can be performed and we find
@@
×$\times\prod_{2\leq k\leq r\leq N}\frac{(1-t_{1k,1r})(1-t_{2k,2r})}{f_{k,r}}\,\prod_{A_{1}=1}^{f}\prod_{1\leq i_{1},j_{1}\leq N}{1\over(1-w_{12}^{A_{1}}{t_{11,1i_{1}}\over t_{21,2j_{1}}})}$
×$\times\prod_{A_{2}=1}^{f}\prod_{1\leq i_{2},j_{2}\leq N}{1\over(1-w_{21}^{A_{2}}{t_{21,2i_{2}}\over t_{11,1j_{2}}})}$

### v2

@@ 1 Introduction
Our focus in this paper is on quantum mechanical systems. There are a number of interesting quantum mechanical systems that have gravity duals. By considering $S$-wave and related truncations of ${\cal N}=4$ super Yang-Mills theory Kim:2003rza ; Okuyama:2002zn ; Ishiki:2006rt ; Harmark:2014mpa , one is able to isolate holographic sectors with a clear bulk dual. These truncations lead to quantum mechanical systems. Finite-$N$ effects also play a central role in the description of black hole microstates, particularly in the discovery of the fortuity mechanism Chang:2022mjp ; Choi:2022caq ; Choi:2023znd ; Chang:2023zqk ; Choi:2023vdm ; Chang:2024zqi : the microstates of $\frac{1}{16}$-BPS black holes fall into two classes, monotone and fortuitous. Monotone states remain BPS at all $N$, while fortuitous states lose their BPS nature above a critical $N$. In the latter case, BPS saturation depends crucially on trace relations. The relevant cohomologies are simpler to study in the BMN matrix model Berenstein:2002jq truncation of the classical field theory.
The number of primary invariants is given by the Krull dimension of the algebra of gauge invariants. For the multi-matrix quantum mechanics of $d$ matrices the Krull dimension is given by $1+(d-1)N^{2}$. In Section 2 we explain how the number of primary invariants can be computed by performing a complete gauge fixing of the theory and counting the number of degrees of freedom that remain. We explain how to reproduce the Krull dimension for multi-matrix models and then go onto to predict the number of primary invariants in a theory of bifundamental fields. In a theory with $f$ bifundamental fields we find that the number of primary invariants is given by
@@ 2 Counting Primary Invariants
$\text{Configuration space}\big/\text{Symmetry group}$ $\equiv$ ${\cal X}\big/G\,$, (3)
@@ 2.1 Multiple Matrices
Notice that all the induced phases depend only on differences $\phi_{a}-\phi_{b}$. Consequently, a common shift $\phi_{a}\to\phi_{a}+\alpha$ leaves all matrices invariant. Hence, of the $N$ residual phases $\phi_{a}$, only N1N-1 are physically meaningful and can be used to further reduce the degrees of freedom.
Since $X^{2}$ is Hermitian, its matrix elements satisfy $m_{ab}=m_{ba}^{*}$. The off-diagonal entries thus contain non-trivial phases. We can use the N1N-1 phase differences $\phi_{a}-\phi_{a+1}$, for $a=1,2,\dots,N-1$, to eliminate the phases of $m_{a\,a+1}$, setting each to be real. These are the additional fields we can fix using the residual symmetry.
We now count the remaining degrees of freedom. There are $(d-1)N^{2}$ matrix elements in the fields $X^{a}$ for $a=2,\dots,d$, and $N$ eigenvalues in the diagonal matrix $X^{1}$. After fixing N1N-1 phases in $X^{2}$ as described above, we find
@@
independent parameters. This correctly counts the number of primary invariants derived from a system of $d>1$ $N\times N$ Hermitian matrices, and agrees with computation of the Krull dimension performed in herstein . This result is intuitively appealing: since the $X^{a}$ are in the adjoint representation, they are invariant under the $U(1)$ in $U(N)\simeq SU(N)\times U(1)$, so that the effective gauge group is $SU(N)$ of dimension $N^{2}-1$. Thus, gauge fixing eliminates $N^{2}-1$ degrees of freedom from the total $dN^{2}$ to leave (8).
@@
matching the number of eigenvalues.
@@ 2.2 Bifundamental fields
$S=\int dt\left(\mbox{Tr\,}\left(\partial_{t}\Gamma^{A}\partial_{t}\bar{\Gamma}^{A}\right)-\mbox{Tr\,}\left(\Gamma^{A}\bar{\Gamma}^{A}\right)\right),$ (11)
@@
that preserve the diagonal and real form of $D$. These $N$ phases are subject to an overall $U(1)$ redundancy (a common shift), so we can use N1N-1 residual parameters to further simplify the remaining fields.
In particular, we apply these residual transformations to $\Gamma^{2}$, which remains unfixed. Since $\Gamma^{2}$ is a complex matrix, the phases in its off-diagonal entries are complex. As in the case of a single Hermitian matrix, we can use N1N-1 relative phase rotations to eliminate the imaginary parts of N1N-1 off-diagonal entries. For example, we may choose to set the phases of the entries $(\Gamma^{2})_{a,a+1}$ to zero for $a=1,\dots,N-1$.
@@ 3 Counting of Secondary Invariants
[…]
@@
The eigenstates are given by polynomials in the creation operators of a fixed degree. We can refine this further by grading with respect to degree in $A^{1\,\dagger}$ and degree in $A^{2\,\dagger}$ separately. Any multi-trace operator, constructed from $n$ $A^{1\,\dagger}$ fields and $m$ $A^{2\,\dagger}$ fields can be written as
@@
where $\delta_{[\sigma]_{r}[\tau]_{r}}$ is 1 if $\sigma]_{r}$ and $\tau]_{r}$ belong to the same restricted conjugacy class and it is zero otherwise. If two permutations $\rho_{1}$ and $\rho_{2}$ are related as
@@
We have indicated where our interest is: in the coefficient $c^{T}_{\alpha N^{2}}$ of the term $x^{\alpha N^{2}}$ which counts the number of operators of degree $\alpha N^{2}$. Since this counts the total contribution at degree $\alpha N^{2}$, the superscript “TT” is for total. We have in mind that $\alpha$ is held fixed as we take $N$ large. This formula is useful because the asymptotic behaviour of the Littlewood-Richardson number needed to estimate the behaviour of the coefficient of $x^{q}$ with $q$ of order $N^{2}$ is known Stanley1 ; Stanley2 ; PakPanova . If $R$ has $n$ boxes, $r$ has $k$ boxes and $s$ has nkn-k boxes, when $n$, $k$ and nkn-k are all large, the largest Littlewood-Richardson coefficient is given by
@@
The result (33) may be modified by finite $N$ constraints. This implies that the Young diagram $R$ must not have more than $N$ rows, It is therefore necessary to inquire about the shape of the Young diagrams $R$, $r$ and $s$ that attain the maximum Littleoowd-Richardson numbers. It turns out, that all three Young diagrams must take the Vershik-Kerov-Logan-Shepp (VKLS) shapeVK1 ; VK2 ; LS . These Young diagrams have a triangular shape and are not cut off by finite $N$ effects for Young diagrams with $\alpha N^{2}$ boxes, as long as $\alpha$ is not too large. Young diagrams with even more boxes take on the truncated VKLS shape LS .
@@
$d_{\alpha N^{2}}<\hat{d}_{\alpha N^{2}}$ $=$ $e^{((\alpha+1)\log(\alpha+1)-\alpha\log(\alpha))N^{2}}$ (48)
@@
From the partition function
$Z(x)$ $=$ $\sum_{q}d_{q}x^{q}\,\Big(1+\sum_{i}c_{i}^{s}x^{i}\Big)$ (55)
we easily obtain the following formula for $c^{T}_{\alpha N^{2}}$
$c_{\alpha N^{2}}^{T}$ $=$ $d_{\alpha N^{2}}+\sum_{i=1}^{\alpha N^{2}}c_{i}^{s}d_{\alpha N^{2}-i}$ (56)
Now, we use the fact that the coefficients $d_{l}$ are strictly increasing with $l$ which follows immediately because the number of ways to partition a positive integer grows as the integer grows [^4] We have two primary invariants of length 1. Thus, the partitions of $l+1$ include (at least) the partitions of $l$ multiplied by the length 1 primaries. Thus $d_{l+1}>d_{l}$.. Consequently we find
$1+\sum_{i=1}^{\alpha N^{2}}c_{i}^{s}$ $\geq$ $1+\sum_{i=1}^{\alpha N^{2}}c_{i}^{s}{d_{\alpha N^{2}-i}\over d_{\alpha N^{2}}}$ (57)
$=$ ${1\over d_{\alpha N^{2}}}\left(d_{\alpha N^{2}}+\sum_{i=1}^{\alpha N^{2}}c_{i}^{s}d_{\alpha N^{2}-i}\right)$ (59)
$=$ ${c^{T}_{\alpha N^{2}}\over d_{\alpha N^{2}}}>{c^{T}_{\alpha N^{2}}\over\hat{d}_{\alpha N^{2}}}$ (61)
$>$ $e^{\alpha\log 2\,N^{2}-{1\over k}\left((k+\alpha)\log(1+{\alpha\over k})-\alpha\log{\alpha\over k}\right)N^{2}}$ (63)
Consequently, we obtain the cumulative lower bound
$1+\sum_{i=1}^{\alpha N^{2}}c_{i}^{s}$ $>$ $e^{\Big(\alpha\log 2-{1\over k}\left((k+\alpha)\log(1+{\alpha\over k})-\alpha\log{\alpha\over k}\right)\Big)N^{2}}$ (64)
To establish our central claim we need to study the sign of the coefficient of $N^{2}$ in the exponent and verify that it is positive.
@@
$\alpha\log 2>{1\over 15}\left((15+\alpha)\log(1+{\alpha\over 15})-\alpha\log{\alpha\over 15}\right)$ (65)
so that the number of secondary invariants must grow as
$1+\sum_{i}c^{s}_{i}$ $>$ $e^{cN^{2}}$ (66)
with $c$ a positive $O(1)$ number, which is the result we wanted to prove.
@@ Acknowledgements
The work of RdMK is supported by a start up research fund of Huzhou University, a Zhejiang Province talent award and by a Changjiang Scholar award. The work of MK is supported by NRF grants RS-2025-25414114, RS-2023-NR077094, RS-2020-NR049598(CQUeST). HJRvZ would like to acknowledge support from the “Quantum Technologies for Sustainable Development” grant from the National Institute for Theoretical and Computational Sciences of South Africa (NITHECS).
@@ Appendix A Molien-Weyl Functions
$\qquad\times\#(\{n_{A_{1}},n_{A_{2}},p_{I_{1}},p_{I_{2}}\})$
where $(\{n_{A_{1}},n_{A_{2}},p_{I_{1}},p_{I_{2}}\})$ is the number of singlets that can be constructed using $n_{A_{1}}$ $\Gamma^{A_{1}}_{12}$ bifundamental fields, $n_{A_{2}}$ $\Gamma^{A_{2}}_{12}$ bifundamental fields, $p_{I_{1}}$ adjoints of the first $U(N)$ factor and $p_{i_{2}}$ adjoints of the second $U(N)$ factor. Using characters we can write the number of singlets as an integral over $U(N)\times U(N)$ to obtain
@@
$\times\,\sum_{n_{A_{1}}=0}^{\infty}(w^{A_{2}}_{21})^{n_{A_{2}}}\chi_{{\rm sym}^{n_{A_{2}}}(F)}(U_{2})\chi_{{\rm sym}^{n_{A_{2}}}(\bar{F})}(U_{1})\prod_{I_{1}=1}^{b_{1}}\sum_{p_{I_{1}}=0}^{\infty}(z_{1}^{I_{1}})^{p_{I_{1}}}\chi_{{\rm sym}^{p_{I_{1}}}(adj)}(U_{1})$
$\qquad\times\prod_{I_{2}=1}^{b_{2}}\sum_{p_{I_{2}}=0}^{\infty}(z_{2}^{I_{2}})^{p_{I_{2}}}\chi_{{\rm sym}^{p_{I_{2}}}(adj)}(U_{2})$
adjadj stands for the adjoint representation. As always, symmetric products of representations appear above because the fields we are using are bosonic. A useful identity is
@@
$\times{\rm exp}\left(\sum_{A_{1}=1}^{f}\sum_{m=1}^{\infty}{(w_{12}^{A_{1}})^{m}\over m}\chi_{F}(U_{1}^{m})\chi_{\bar{F}}(U_{2}^{m})+\sum_{A_{2}=1}^{f}\sum_{m=1}^{\infty}{(w_{21}^{A_{2}})^{m}\over m}\chi_{F}(U_{2}^{m})\chi_{\bar{F}}(U_{1}^{m})\right)$
@@
$\times\prod_{I_{1}=1}^{b_{1}}{1\over(1-z_{1}^{I_{1}})^{N}}\prod_{1\leq k<r\leq N}{1\over(1-{\varepsilon_{1r}\over\varepsilon_{1k}}z_{1}^{I_{1}})(1-{\varepsilon_{1k}\over\varepsilon_{1r}}z_{1}^{I_{1}})}$
$\times\prod_{I_{2}=1}^{b_{2}}{1\over(1-z_{2}^{I_{2}})^{N}}\prod_{1\leq s<t\leq N}{1\over(1-{\varepsilon_{2t}\over\varepsilon_{2s}}z_{2}^{I_{2}})(1-{\varepsilon_{2s}\over\varepsilon_{2t}}z_{2}^{I_{2}})}$
$\times\prod_{A_{1}=1}^{f}\prod_{1\leq i_{1},j_{1}\leq N}{1\over(1-w_{12}^{A_{1}}{\varepsilon_{1i_{1}}\over\varepsilon_{2j_{1}}})}\prod_{A_{2}=1}^{f}\prod_{1\leq i_{2},j_{2}\leq N}{1\over(1-w_{21}^{A_{2}}{\varepsilon_{2i_{2}}\over\varepsilon_{1j_{2}}})}$
We now simplify the integrand of the above formula. According to (90) and (91) each of the products $\Delta\bar{\Delta}$ involves a sum over permutations $\sigma$ and $\tau$. For each distinct term in the sum over $\sigma$, consider the distinct change of variables given by a permutation of the eigenvalues $\varepsilon_{1,i}\to\varepsilon_{1,\sigma(i)}$, $\varepsilon_{2,i}\to\varepsilon_{2,\sigma(i)}$. All factors in the integrand are invariant under this change except for the Van der Monde determinants which are both either invariant or both pick up a minus sign. Consequently each product $\Delta\bar{\Delta}$ is invariant. The net result of this change of variables is that the sum over $\sigma$ can be performed and we find
@@
$\times\prod_{2\leq k\leq r\leq N}\frac{(1-t_{1k,1r})(1-t_{2k,2r})}{f_{k,r}}\,\prod_{A_{1}=1}^{f}\prod_{1\leq i_{1},j_{1}\leq N}{1\over(1-w_{12}^{A_{1}}{t_{11,1i_{1}}\over t_{21,2j_{1}}})}$
$\times\prod_{A_{2}=1}^{f}\prod_{1\leq i_{2},j_{2}\leq N}{1\over(1-w_{21}^{A_{2}}{t_{21,2i_{2}}\over t_{11,1j_{2}}})}$